OpenStudy (anonymous):

What is the additive inverse of the complex number 9 – 4i?

zepdrix (zepdrix):

Additive inverse should get you to zero when you apply it:\[\large\rm 9-4i+z=0\]

zepdrix (zepdrix):

Where z is some complex number, x+iy, \[\large\rm 9-4i+x+iy=0\]Equate the real parts,\[\large\rm 9+x=0\]Solve for x, and equate the imaginary parts,\[\large\rm -4i+yi=0\]and solve for y.

OpenStudy (anonymous):

so –9 + 4i

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

what about What is the value of the product (3 – 2i)(3 + 2i)?

OpenStudy (anonymous):

5 9 + 4i 9 – 4i 13

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

You can FOIL it out, \(\rm (a+bi)(a-bi)=a^2+abi-abi-b^2i^2\) But what you'll find is that the middle terms always cancel out, and what you're left with, \(\rm =a^2+b^2\) is the sum of squares. This only applies when the brackets are exactly the same, but with a different sign between the terms. we call these conjugates. So therefore, by that rule: \(\rm (3-2i)(3+2i)=3^2+2^2\)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

3^2+2^2 is 5? Hmm

OpenStudy (anonymous):

9 + 4i